OpenStudy (anonymous):

What is the solution to the system of equations?' x+y=2 x-y=4

OpenStudy (anonymous):

@mathstudent55

OpenStudy (anonymous):

@Straybullet

OpenStudy (anonymous):

@texaschic101

OpenStudy (anonymous):

@Darknight200

OpenStudy (texaschic101):

x + y = 2 x - y = 4 ---------add 2x = 6 -- divide both sides by 2 x = 3 now sub 3 in for x in either equation to find y x + y = 2 3 + y = 2 y = 2 - 3 y = -1 so x = 3 and y = -1 any questions ?

OpenStudy (anonymous):

1%2Ax%2B1%2Ay=2 1%2Ax-1%2Ay=4 In order to solve for one variable, we must eliminate the other variable. So if we wanted to solve for y, we would have to eliminate x (or vice versa). So lets eliminate x. In order to do that, we need to have both x coefficients that are equal but have opposite signs (for instance 2 and -2 are equal but have opposite signs). This way they will add to zero. So to make the x coefficients equal but opposite, we need to multiply both x coefficients by some number to get them to an equal number. So if we wanted to get 1 and 1 to some equal number, we could try to get them to the LCM. Since the LCM of 1 and 1 is 1, we need to multiply both sides of the top equation by 1 and multiply both sides of the bottom equation by -1 like this: 1%2A%281%2Ax%2B1%2Ay%29=%282%29%2A1 Multiply the top equation (both sides) by 1 -1%2A%281%2Ax-1%2Ay%29=%284%29%2A-1 Multiply the bottom equation (both sides) by -1 So after multiplying we get this: 1%2Ax%2B1%2Ay=2 -1%2Ax%2B1%2Ay=-4 Notice how 1 and -1 add to zero (ie 1%2B-1=0) Now add the equations together. In order to add 2 equations, group like terms and combine them %281%2Ax-1%2Ax%29%2B%281%2Ay%2B1%2Ay%29=2-4 %281-1%29%2Ax%2B%281%2B1%29y=2-4 cross%281%2B-1%29%2Ax%2B%281%2B1%29%2Ay=2-4 Notice the x coefficients add to zero and cancel out. This means we've eliminated x altogether. So after adding and canceling out the x terms we're left with: 2%2Ay=-2 y=-2%2F2 Divide both sides by 2 to solve for y y=-1 Reduce Now plug this answer into the top equation 1%2Ax%2B1%2Ay=2 to solve for x 1%2Ax%2B1%28-1%29=2 Plug in y=-1 1%2Ax-1=2 Multiply 1%2Ax=2%2B1 Subtract -1 from both sides 1%2Ax=3 Combine the terms on the right side cross%28%281%2F1%29%281%29%29%2Ax=%283%29%281%2F1%29 Multiply both sides by 1%2F1. This will cancel out 1 on the left side. x=3 Multiply the terms on the right side So our answer is x=3, y=-1 which also looks like (3, -1) Notice if we graph the equations (if you need help with graphing, check out this solver) 1%2Ax%2B1%2Ay=2 1%2Ax-1%2Ay=4 we get graph of 1%2Ax%2B1%2Ay=2 (red) 1%2Ax-1%2Ay=4 (green) (hint: you may have to solve for y to graph these) and the intersection of the lines (blue circle). and we can see that the two equations intersect at (3,-1). This verifies our answer.

OpenStudy (anonymous):

I think its x=3, y=-1 tell me if im wrong